# Scalable Quantum Simulations of Scattering in Scalar Field Theory on 120 Qubits

### v1

@@ Appendix C Error mitigation
Instead, the circuits simulating the vacuum time evolution are used as mitigation circuits in this work. If the vacuum were prepared exactly, and if the time evolution operator were implemented exactly, time evolution would have no effect as the vacuum is an eigenstate of the Hamiltonian, $e^{-itH}|\psi_{\text{vac}}\rangle=|\psi_{\text{vac}}\rangle$. Because both the vacuum state preparation and the time evolution are implemented approximately using variational circuits, the prepared $|\psi_{\text{vac}}\rangle$ has some time evolution. However, this vacuum evolution can be computed classically, owing to the translation-invariant nature of $|\psi_{\text{vac}}\rangle$. Up to finite-size effects, the time evolution of translationally-invariant states does not change with system size. The vacuum time evolution for a given system size can be estimated following the same procedure as for the convergence of the variational parameters described in the SVC algorithm. Due to the fact that correlations in the vacuum decay exponentially in this theory, finite-size effects have an exponentially-small contribution. As a result, the infinite-volume vacuum expectation values $\langle\psi_{\text{vac}}|ZZ|\psi_{\text{vac}}\rangle_{\infty}$ (as well as those for any $L$) can be estimated by extrapolating from an exponential fit to the vacuum expectation values at a series of $L$. This provides a way to compute the $\langle ZZ\rangle_{\text{mit,true}}$. The values of these extrapolations for all simulation times in both the free and interacting theories, as well as the values at $L=60$ determined by MPS, are provided in Table 4. For the parameters and simulation times considered, this method is seen to give much better estimates of the noise, as a result of the same time evolution being applied in both the physics and mitigation circuits. The effect of ODR can be seen in Fig. 13. The unmitigated data shown in the upper panel is much closer to the completely decohered result (grey dotted line), than the true result from MPS (grey dashed line). After the error mitigation scheme is applied (bottom panel) to measure the noise and adjust the results, measurements of $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}$ are much closer to the expected results from MPS simulations. The action of ODR on $\langle ZZ\rangle_{\text{2wp}}$ measurements, from which $\langle\phi^{2}_{j}\rangle_{\text{2wp}}$ is calculated, is shown in the insets of Fig. 13 for lattice site $j=4$.
@@
$t$ Even sites Odd sites Even sites Odd sites
@@
Table 4: $\langle\psi_{\text{vac}}|e^{itH}\phi^{2}_{j}e^{-itH}|\psi_{\text{vac}}\rangle$ that are used as “true” values in the error mitigation scheme. Expectation values determined from classical simulations by using an exponential extrapolation are compared to those determined by MPS. A maximum system size of $L=16$ (32 qubits) was used to compute the extrapolation. The even and odd sites have slightly different values because of the Trotterized state preparation, see Sec. IV.1 for details.
Figure 13: The effect of ODR and filtering on the data from ibm_fez for $\lambda=0,\,t=8$. The points show data collected using 80 PTs, two TREX twirls, and 8000 shots for each circuit. The expectation value $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}$ is shown before (top panel) and after (bottom panel) ODR and filtering. The grey dashed line represents classical noiseless MPS simulations, and the grey dotted line shows the results if the state had fully decohered to the completely mixed state. The error bars represent one standard deviation of the bootstrap-resampled data. ODR and filtering act on the $\langle ZZ\rangle_{\text{2wp}}$ expectation value twirl by twirl (insets).
@@
Note that it is possible to directly calculate an approximation to the vacuum evolution using MPS. In this work, results from classical MPS simulations are used as ground truth to compare to results from ibm_fez. For the system parameters chosen and the times considered, it can be seen that the extrapolated vacuum evolution is quite close to the MPS predictions (see Table 4), with the differences only noticeable at late times. It is expected that highly-inelastic collisions will generate large quantities of entanglement and magic, thus being genuinely out of reach for simulations using classical methods such as MPS or Clifford+T. In this setting, the MPS values for the vacuum evolution may be used directly for mitigation. This is a reasonable expectation, as the ground states of one-dimensional gapped systems, such as the scalar field theory considered in this work, are known to obey area law entanglement [187], so their short-time evolution may be simulated with an MPS with a relatively small bond dimension. [^21] The approximate variational methods may yield states that are more entangled than the true, non-variational states. This may be averted by using techniques similar to Refs. [144, 104, 147] to train the variational circuits with MPS in the first place. Furthermore, MPS approximations to the $|\psi_{\text{2wp}}\rangle$ evolution circuit may be used as mitigation circuits in the future as well, provided they can be constructed in a way that preserves the structure of the noise.
@@
The IBM Heron processors have seen a 3-5$\times$ increase in performance over their Eagle predecessors, and have significantly decreased qubit crosstalk [188]. Despite this, it is still seen in this work that crosstalk remains a non-negligible source of error. DD is used to minimize this effect and remove erroneous evolution during qubit idle times. To mitigate these effects, the crosstalk-robust CRXY4 DD sequence [189] is used. This sequence consists of staggered XY4 sequences applied to neighboring qubits to cancel out collective evolution caused by crosstalk. [^22] Recent work found that asymmetric implementations of $Y$ gates can have an adverse impact on the effectiveness of DD sequences [190]. As a result, all $Y$ gates used in this work, including the DD stage, are converted to the native gate set using the symmetric implementation $Y=R_{Z}(\pi/2)XR_{Z}(-\pi/2)$.
@@
Appendix D Variational parameters
In this section, the variational parameters determined using the SVC framework presented in Sec. III are given for vacuum preparation, wavepacket excitation, and time evolution. Exponential extrapolation is run for the vacuum preparation parameters. The $L$-extrapolated vacuum parameters are used at $L=12$ to optimize the wavepacket preparation and time evolution circuits.
@@ Appendix D Variational parameters
Table 6: The parameters $\theta_{ij}$ used to initialize Gaussian wavepackets of momenta $k=\pm\pi/3$ and spread in $k$-space $\sigma_{k}=\pi/3$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3c: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
@@
11 1.8584 1.1814 1.594 -0.0322 0.3009 -0.0683 1.6193 -1.1086 0.9842 -0.5982 -0.0874 0.0748
Table 7: The parameters $\theta_{ij}$ used to implement time evolution for $t=1$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
@@
Table 8: The parameters $\theta_{ij}$ used to implement time evolution for $t=2$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
@@
11 3.1055 1.5655 1.6030 0.2539 -0.0241 -0.0084 1.0308 0.1463 1.670 -0.0398 -1.8732 -0.0112
Table 9: The parameters $\theta_{ij}$ used to implement time evolution for $t=3$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
@@
Table 10: The parameters $\theta_{ij}$ used to implement time evolution for $t=4$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
@@
11 3.9799 1.1996 1.6125 0.2875 -1.0267 -0.0509 -2.6706 -0.0246 1.5695 -$0$1.5884 0.1631
Table 11: The parameters $\theta_{ij}$ used to implement time evolution for $t=5$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
@@
Table 12: The parameters $\theta_{ij}$ used to implement time evolution for $t=6$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
@@
Table 13: The parameters $\theta_{ij}$ used to implement time evolution for $t=7$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
@@
Table 14: The parameters $\theta_{ij}$ used to implement time evolution for $t=8$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
@@
Table 15: The parameters $\theta_{ij}$ used to implement time evolution for $t=9$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer $i$. These parameters are determined using classical simulations and optimization.
Appendix E Tables
In this section, expectation values $\langle\phi_{j}^{2}\rangle_{\text{2wp}}-\langle\phi_{j}^{2}\rangle_{\text{vac}}$ are tabulated for $\lambda=0$ and $\lambda=2$. Error mitigated results from ibm_fez are compared to classical MPS data. Larger uncertainties in the data from the quantum computer can be seen as a result of increasing circuit depth: the circuits for $t=1,2,3$ have a depth of 59 (Tables 16, 17, 18), those for $t=4,5,6$ have a depth of 81 (Tables 19, 20, 21), and those for $t=7,8,9$ have a depth of 103 (Tables 22, 23, 24).
@@ Appendix E Tables
Spatial site $j$ $\lambda=0$ $\lambda=2$
@@
Spatial site $j$ $\lambda=0$ $\lambda=2$
@@
Spatial site $j$ $\lambda=0$ $\lambda=2$
@@
Spatial site $j$ $\lambda=0$ $\lambda=2$
@@
Spatial site $j$ $\lambda=0$ $\lambda=2$
@@
Spatial site $j$ $\lambda=0$ $\lambda=2$
@@
Spatial site $j$ $\lambda=0$ $\lambda=2$
@@
Spatial site $j$ $\lambda=0$ $\lambda=2$
@@
Spatial site $j$ $\lambda=0$ $\lambda=2$

### v2

@@ Appendix E Error mitigation
Instead, the circuits simulating the vacuum time evolution are used as mitigation circuits in this work. If the vacuum were prepared exactly, and if the time evolution operator were implemented exactly, time evolution would have no effect as the vacuum is an eigenstate of the Hamiltonian, $e^{-itH}|\psi_{\text{vac}}\rangle=|\psi_{\text{vac}}\rangle$. Since both the vacuum state preparation and the time evolution are implemented approximately using variational circuits, the prepared $|\psi_{\text{vac}}\rangle$ has some time evolution. However, this vacuum evolution can be computed classically, owing to the translation-invariant nature of $|\psi_{\text{vac}}\rangle$. Up to finite-size effects, the time evolution of translationally-invariant states does not change with system size. The vacuum time evolution for a given system size can be estimated following the same procedure as for the convergence of the variational parameters described in the SVC algorithm. Due to the fact that correlations in the vacuum decay exponentially in this theory, finite-size effects have an exponentially-small contribution. As a result, the infinite-volume vacuum expectation values $\langle\psi_{\text{vac}}|ZZ|\psi_{\text{vac}}\rangle_{\infty}$ (as well as those for any $L$) can be estimated by extrapolating from an exponential fit to the vacuum expectation values at a series of $L$. This provides a way to compute the $\langle ZZ\rangle_{\text{mit,true}}$. The values of these extrapolations for all simulation times in both the free and interacting theories, as well as the values at $L=60$ determined by MPS, are provided in Table 5. For the parameters and simulation times considered, this method is seen to give much better estimates of the noise, as a result of the same time evolution being applied in both the physics and mitigation circuits. The effect of ODR can be seen in Fig. 15. The unmitigated data shown in the upper panel is much closer to the completely decohered result (grey dotted line), than the true result from MPS circuit simulations (grey dashed line). After the error mitigation scheme is applied (bottom panel) to measure the noise and adjust the results, measurements of $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}$ are much closer to the expected results from MPS. The action of ODR on $\langle ZZ\rangle_{\text{2wp}}$ measurements, from which $\langle\phi^{2}_{j}\rangle_{\text{2wp}}$ is calculated, is shown in the insets of Fig. 15 for lattice site $j=4$.
@@
$t$
Even sites Odd sites Even sites Odd sites
@@
Table 5: $\langle\psi_{\text{vac}}|e^{itH}\phi^{2}_{j}e^{-itH}|\psi_{\text{vac}}\rangle$ that are used as “true” values in the error mitigation scheme for the $L=60$ (120 qubits) simulations. Expectation values determined from classical statevector simulations by using an exponential extrapolation are compared to those determined by a MPS circuit simulation of the $L=60$ system. A maximum system size of $L=16$ (32 qubits) was used to compute the extrapolation. The even and odd sites have slightly different values because of the Trotterized state preparation, see Sec. IV.1 for details.
Figure 15: The effect of ODR and filtering on the data from ibm_fez for $\lambda=0,\,t=8$. The points show data collected using 80 PTs, two TREX twirls, and 8000 shots for each circuit. The expectation value $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}$ is shown before (top panel) and after (bottom panel) ODR and filtering. The grey dashed line represents classical noiseless MPS circuit simulations, and the grey dotted line shows the results if the state had fully decohered to the completely mixed state. The error bars represent one standard deviation of the bootstrap-resampled data. ODR and filtering act on the $\langle ZZ\rangle_{\text{2wp}}$ expectation value twirl by twirl (insets).
@@
Note that it is possible to directly calculate an approximation to the vacuum evolution using MPS. In this work, results from classical MPS circuit simulations are used as ground truth to compare to results from ibm_fez. For the system parameters chosen and the times considered, it can be seen that the extrapolated vacuum evolution is quite close to the MPS predictions (see Table 5), with the differences only noticeable at late times. It is expected that highly-inelastic collisions will generate large quantities of entanglement and magic, thus being genuinely out of reach for simulations using classical methods such as MPS or Clifford+T. In this setting, the MPS values for the vacuum evolution may be used directly for mitigation. This is a reasonable expectation, as the ground states of one-dimensional gapped systems, such as the scalar field theory considered in this work, are known to obey area law entanglement [196], so their short-time evolution may be simulated with an MPS with a relatively small bond dimension. [^25] The approximate variational methods may yield states that are more entangled than the true, non-variational states. This may be averted by using techniques similar to Refs. [151, 110, 154] to train the variational circuits with MPS in the first place. Furthermore, MPS approximations to the $|\psi_{\text{2wp}}\rangle$ evolution circuit may be used as mitigation circuits in the future as well, provided they can be constructed in a way that preserves the structure of the noise.
@@
[…]
@@
Appendix F Variational parameters
In this appendix, the variational parameters determined using the SVC framework presented in Sec. III are given for vacuum preparation, wavepacket excitation, and time evolution. Exponential extrapolation is run for the vacuum preparation parameters. The $L$-extrapolated vacuum parameters are used at $L=12$ to optimize the wavepacket preparation and time evolution circuits.
@@ Appendix F Variational parameters
Table 7: The parameters $\theta_{ij}$ used to initialize Gaussian wavepackets of momenta $k=\pm\pi/3$ and spread in $k$-space $\sigma_{k}=\pi/3$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3c: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
@@
11 1.8584 1.1814 1.5940 -0.0322 0.3009 -0.0683 1.6193 -1.1086 0.9842 -0.5982 -0.0874 0.0748
Table 8: The parameters $\theta_{ij}$ used to implement time evolution for $t=1$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
@@
Table 9: The parameters $\theta_{ij}$ used to implement time evolution for $t=2$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
@@
11 3.1055 1.5655 1.6030 0.2539 -0.0241 -0.0084 1.0308 0.1463 1.6700 -0.0398 -1.8732 -0.0112
Table 10: The parameters $\theta_{ij}$ used to implement time evolution for $t=3$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
@@
Table 11: The parameters $\theta_{ij}$ used to implement time evolution for $t=4$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
@@
11 3.9799 1.1996 1.6125 0.2875 -1.0267 -0.0509 -2.6706 -0.0246 1.5695 -0.0200 -1.5884 0.1631
Table 12: The parameters $\theta_{ij}$ used to implement time evolution for $t=5$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
@@
Table 13: The parameters $\theta_{ij}$ used to implement time evolution for $t=6$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
@@
Table 14: The parameters $\theta_{ij}$ used to implement time evolution for $t=7$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
@@
Table 15: The parameters $\theta_{ij}$ used to implement time evolution for $t=8$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
@@
Table 16: The parameters $\theta_{ij}$ used to implement time evolution for $t=9$ in the free and interacting theories. The $\theta_{ij}$ correspond to the angles defined in the brickwall layer shown in Fig. 3d: $i$ labels the layer, and $j$ labels the parameter within layer ii. These parameters are determined using classical simulations and optimization.
Appendix G Tables
In this appendix, expectation values $\langle\phi_{j}^{2}\rangle_{\text{2wp}}-\langle\phi_{j}^{2}\rangle_{\text{vac}}$ are tabulated for $\lambda=0$ and $\lambda=2$. Error-mitigated results from ibm_fez are compared to classical data from a MPS circuit simulator. Larger uncertainties in the data from the quantum computer can be seen as a result of increasing circuit depth: the circuits for $t=1,2,3$ have a depth of 59 (Tables 17, 18, 19), those for $t=4,5,6$ have a depth of 81 (Tables 20, 21, 22), and those for $t=7,8,9$ have a depth of 103 (Tables 23, 24, 25).
@@ Appendix G Tables
Spatial site
$j$
$\lambda=0$ $\lambda=2$
@@
Spatial site
$j$
$\lambda=0$ $\lambda=2$
@@
Spatial site
$j$
$\lambda=0$ $\lambda=2$
@@
Spatial site
$j$
$\lambda=0$ $\lambda=2$
@@
Spatial site
$j$
$\lambda=0$ $\lambda=2$
@@
Spatial site
$j$
$\lambda=0$ $\lambda=2$
@@
Spatial site
$j$
$\lambda=0$ $\lambda=2$
@@
Spatial site
$j$
$\lambda=0$ $\lambda=2$
@@
Spatial site
$j$
$\lambda=0$ $\lambda=2$